# Beyond Linear Approximations: A Novel Pruning Approach for Attention Matrix

### v1

@@
Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Jiangxuan Long lungchianghsuan@gmail.com. South China University of Technology.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song magic.linuxkde@gmail.com. The Simons Institute for the Theory of Computing at the University of California, Berkeley.    Yufa Zhou yufazhou@seas.upenn.edu. University of Pennsylvania.
Large Language Models (LLMs) have shown immense potential in enhancing various aspects of our daily lives, from conversational AI to search and AI assistants. However, their growing capabilities come at the cost of extremely large model sizes, making deployment on edge devices challenging due to memory and computational constraints. This paper introduces a novel approach to LLM weight pruning that directly optimizes for approximating the attention matrix, a core component of transformer architectures. Unlike existing methods that focus on linear approximations, our approach accounts for the non-linear nature of the Softmax attention mechanism. We provide theoretical guarantees for the convergence of our Gradient Descent-based optimization method to a near-optimal pruning mask solution. Our preliminary empirical results demonstrate the effectiveness of this approach in maintaining model performance while significantly reducing computational costs. This work establishes a new theoretical foundation for pruning algorithm design in LLMs, potentially paving the way for more efficient LLM inference on resource-constrained devices.
@@ 1 Introduction
Large Language Models (LLMs) based on the transformer architecture [81], including GPT-4o [68], Claude [5], and OpenAI’s recent o1 [69], have shown immense potential to enhance our daily lives. They revolutionize fields like conversational AI [51], AI agents [85, 17], search AI [69], and AI assistants [66, 94, 47, 30]. With their growing capabilities, LLMs are powerful tools shaping the future of technology. However, the current state-of-the-art LLM weights number is extremely large. For instance, the smallest version of Llama 3.1 [62] needs 8 billion parameters, which takes more than 16GB GPU memory with half float precision and requires significant inference time. Due to large memory and high computational cost, deploying such models on edge devices such as smartphones becomes challenging.
To reduce the LLM model size, many studies work on pruning the LLMs model weights to relax the device memory constraint and minimize response latency. The classical pruning problem in LLMs can be formulated as follows. Given a weight matrix $W\in\mathbb{R}^{d\times d}$ and some calibration data $X\in\mathbb{R}^{n\times d}$, where $n$ is input token length and $d$ is feature dimension, the goal is to find a matrix $\widetilde{W}$ under some sparse constraint such that $\|XW-X\widetilde{W}\|$ being small under some norm function. The above formulation has been widely used in many state-of-the-art pruning methods, such as SparseGPT [26] and Wanda [75].
However, the current object functions only focus on the approximation of a linear function $XW$. Their optimal solutions do not have a good approximation to the attention matrix (see Figure 2 for details). Note that the attention mechanism is the kernel module of the transformer architecture. The high-level insight of their bad performance is that the $\mathsf{Softmax}$ function is very sensitive to the large positive values of the input due to its $\exp$ scaling effect while pruning mask based on linear approximation cannot capture this sensitivity. Thus, in this work, we directly compute the pruning mask on weights to approximate the attention matrix, which is a highly non-linear function, $\mathsf{Softmax}(XWX^{\top})\in\mathbb{R}^{n\times n}$. To the best of our knowledge, this paper is the first work studying attention weight pruning to directly approximate the attention matrix. We provide a theoretical guarantee that optimization based on Gradient Descent (GD) on our loss function can converge to a good pruning mask solution (Theorem 1.3). Furthermore, we preliminarily verified the effectiveness of our method with empirical support (Section 6). Our theoretical foundation may pave the way for more efficient LLM inference on resource-constrained devices.
In the following, we introduce some key backgrounds and our contributions in detail.
@@
1.1 Key Background
We define the attention matrix in self-attention mechanism as below:
@@ 1.1 Key Background
Let $M\in[0,1]^{d\times d}$ be the pruning mask. Let $X,W$ be defined in Definition 1.1. Let $A:=\exp(XWX^{\top})$ and $\widetilde{A}:=\exp(X(M\circ W)X^{\top})$, where $\circ$ is the Hadamard product. Let $D:=\operatorname{diag}(A\cdot{\bf 1}_{n})$ and $\widetilde{D}:=\operatorname{diag}(\widetilde{A}\cdot{\bf 1}_{n})$. Let $\lambda\in\mathbb{R}_{+}$ be the regularization parameter. We define Attention Weights Pruning loss function to be
@@
1.2 Our Contributions
This is the first work studying the Attention Weights Pruning problem, which is an approximation problem to a non-linear function. We provide an algorithm to obtain the near-optimal pruning mask based on Gradient Descent (GD) with convergence guarantee.
@@ 1.2 Our Contributions
In the above theorem, $\xi$ can be arbitrarily small as $\xi\rightarrow 0$ when the regularization coefficient $\lambda\rightarrow 0$. So our analysis shows that although the objective function is highly non-linear, the GD training can converge to a near-optimal pruning mask solution, supported by our experiments in Section 6.
@@
  • We conduct preliminary experiments to verify the effectiveness of our method (Section 6).
@@
Our paper is organized as follows. In Section 2, we review the related work. Section 3 introduces key concepts and definitions essential for the subsequent sections. In Section 4, we present our main result. Section 5 offers a technical overview of the methods employed. Experimental results are discussed in Section 6. Finally, Section 7 summarizes our findings and offers concluding remarks.
@@ 2.1 Pruning and Compression for LLMs
Model compression plays a critical role in improving the efficiency and deployment of large language models (LLMs) [96] for its effectiveness in reducing computational overhead while preserving performance. Common compression techniques include quantization [70, 88, 36], pruning [13, 32, 33, 44, 25, 26, 75, 55, 92, 97, 90, 2, 15], and knowledge distillation [38, 77, 43, 84]. Specifically, pruning techniques have been developed extensively, such as unstructured pruning, which removes individual weights [55, 75], and structured pruning, which eliminates entire components like neurons or attention heads [64, 2, 86]. [75] proposed Wanda, a novel unstructured pruning technique that uses weight-activation products to induce up to 50% sparsity in LLMs without retraining, achieving competitive results with significantly lower computational cost. SparseGPT [26] introduced a one-shot pruning method that achieves up to 60% sparsity in large GPT-family models with minimal impact on performance. A follow-up work [55] improved the complexity analysis of SparseGPT, reducing the running time from $O(d^{3})$ to $O(d^{2.53})$, enabling faster pruning on LLMs. These techniques together contribute to more scalable and resource-efficient LLMs, maintaining competitive performance while having substantial reductions in computational resources.
@@ 2.2 Attention Acceleration
Attention mechanism has faced criticism due to its quadratic time complexity with respect to context length [81]. Addressing this criticism, a variety of approaches are employed, including sparse attention [33, 48, 26, 55], low-rank approximations [73, 53, 40, 95, 39], and kernel-based methods [14, 63, 21, 93, 58], to reduce computational overhead and improve scalability. [1] enable the derivation of a low-rank representation of the attention matrix, which accelerates both the training and inference processes of single attention layer, tensor attention, and multi-layer transformer, achieving almost linear time complexity [6, 7, 8, 60, 57]. Other approaches like Mamba [31, 20], Linearizing Transformers [91, 65], Hopfield Models [42, 83, 37, 87, 82, 34, 35, 41], and PolySketchFormer [49] focus on architectural modifications and implementation optimizations to enhance performance. System-level optimizations such as FlashAttention [19, 18, 74] and block-wise parallel decoding [78] further improve efficiency. Collectively, these innovations have significantly augmented transformer models’ ability to handle longer input sequences, unlocking broader applications across multiple sectors [16, 72, 24, 67, 89, 4, 45, 54, 59, 76].
@@ 3 Preliminary
In this section, we introduce some basic concepts and key definitions. In Section 3.1, we introduce some basic notations we use in this paper. In Section 3.2, we provide the definition of attention weights pruning.
3.1 Notations
For any positive integer $n$, we use $[n]$ to denote set $\{1,2,\cdots,n\}$. For each $a,b\in\mathbb{R}^{n}$, we use $a\circ b\in\mathbb{R}^{n}$ to denote the Hadamard product, i.e., the $i$-th entry of $(a\circ b)$ is $a_{i}b_{i}$ for all $i\in[n]$. For $A\in\mathbb{R}^{m\times n}$, let $A_{i}\in\mathbb{R}^{n}$ denote the $i$-th row and $A_{*,j}\in\mathbb{R}^{m}$ denote the $j$-th column of $A$, where $i\in[m]$ and $j\in[n]$. We use $\exp(A)$ to denote a matrix where $\exp(A)_{i,j}:=\exp(A_{i,j})$ for a matrix $A\in\mathbb{R}^{n\times d}$. We use $\|A\|_{F}$ to denote the Frobenius norm of a matrix $A\in\mathbb{R}^{n\times d}$, i.e., $\|A\|_{F}:=\sqrt{\sum_{i\in[n]}\sum_{j\in[d]}|A_{i,j}|^{2}}$. For a symmetric matrix $A\in\mathbb{R}^{n\times n}$, $A\succeq 0$ means that $A$ is positive semidefinite (PSD), i.e., for all $x\in\mathbb{R}^{n}$, we have $x^{\top}Ax\geq 0$.
3.2 Attention Weights Pruning
We aim to determine a near-optimal pruning mask $M$ for the attention weights in a self-attention mechanism. Furthermore, we incorporate causal attention masking [^1] In this paper, we always use pruning mask to refer $M\in\mathbb{R}^{d\times d}$, which is our target, and use causal attention mask to refer $M_{c}\in\mathbb{R}^{n\times n}$, which is a fixed mask for standard self-attention. into our method to be more aligned with the current decoder-only LLM architecture, while our analysis can be applied to any general attention mask, e.g., block-wise attention mask. To formalize this, we provide the formal definition of causal attention mask and attention weights pruning in this section.
The causal attention mask ensures that each token in the sequence can attend only to itself and preceding tokens. Here, we provide the formal definition for the causal attention mask:
Definition 3.1 (Causal attention mask, [57]).
Let $M_{c}\in\{0,1\}^{n\times n}$ be the causal attention mask defined in Definition 3.1. Let $A:=\exp(XWX^{\top})\circ M_{c}$ and $\widetilde{A}:=\exp(X(M\circ W)X^{\top})\circ M_{c}$. Let $D:=\operatorname{diag}(A\cdot{\bf 1}_{n})$ and $\widetilde{D}:=\operatorname{diag}(\widetilde{A}\cdot{\bf 1}_{n})$. We define Attention Weights Pruning with Causal Attention Mask loss function to be
$\mathcal{L}(M):=\mathcal{L}_{\mathrm{attn}}(M)+\mathcal{L}_{\mathrm{reg}}(M)$
where
$\mathcal{L}_{\mathrm{attn}}(M):=\frac{1}{2}\|D^{-1}A-\widetilde{D}^{-1}\widetilde{A}\|_{F}^{2}~{}~{}\mathrm{and}~{}~{}\mathcal{L}_{\mathrm{reg}}(M):=\frac{1}{2}\lambda\|M\|_{F}^{2}.$
@@ 4 Main Results
In this section, we provide our main results. We provide an Algorithm 1 for Attention Weights Pruning problem based on Gradient Descent (GD). We also prove the convergence for our GD algorithm in Theorem 4.1.
@@
Let $M,~{}X,~{}W,\widetilde{D},\widetilde{A}~{}\lambda,~{}\mathcal{L},~{}\mathcal{L}_{\mathrm{attn}}$ be defined in Definition 3.2. Assume $XX^{\top}\succeq\beta I$ and $\min_{i,j\in[n]}(\widetilde{D}^{-1}\widetilde{A})_{i,j}\geq\delta>0$. Furthermore,
  • Let $\mu=2\min_{i,j\in[d]}\{|W_{i,j}|\}\cdot\beta\cdot\delta$.
  • Let $\xi=12n^{-1.5}\max_{i,j\in[d]}\{|W_{i,j}|\}\cdot\|X\|_{F}^{2}\cdot\lambda d/\mu$.
Then, for any $\epsilon>0$, provided $\eta<1/L$ where $L$ is the Lipschitz constant for $\nabla_{M}\mathcal{L}(M)$ (see Theorem 5.4), GD (Algorithm 1) with fixed step size $\eta$ and run for $T=4\mathcal{L}(M^{(0)})/(\eta\mu\epsilon n^{2})$ iterations results in the following guarantee,
@@
After solving the optimization problem, we obtain a pruning mask with real-valued entries. In practice, however, this pruning mask must be converted into a binary form, specifically $M\in\{0,1\}^{d\times d}$. We define the pruning ratio $\rho$ as the percentage of weights to be pruned. We apply this ratio by setting the pruning mask entries to zero for weights that fall below the $\rho$-th percentile and to one for those above. This ensures that only the specified proportion of weights are pruned.

### v2

@@
Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Jiangxuan Long lungchianghsuan@gmail.com. South China University of Technology.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song magic.linuxkde@gmail.com. The Simons Institute for the Theory of Computing at the UC, Berkeley.    Yufa Zhou yufazhou@seas.upenn.edu. University of Pennsylvania.
Large Language Models (LLMs) have shown immense potential in enhancing various aspects of our daily lives, from conversational AI to search and AI assistants. However, their growing capabilities come at the cost of extremely large model sizes, making deployment on edge devices challenging due to memory and computational constraints. This paper introduces a novel approach to LLM weight pruning that directly optimizes for approximating the attention matrix, a core component of transformer architectures. Unlike existing methods that focus on linear approximations, our approach accounts for the non-linear nature of the Softmax attention mechanism. We provide theoretical guarantees for the convergence of our Gradient Descent-based optimization method to a near-optimal pruning mask solution. Our empirical results demonstrate the effectiveness of our non-linear pruning approach in maintaining model performance while significantly reducing computational costs, which is beyond the current state-of-the-art methods, i.e., SparseGPT and Wanda, by a large margin. This work establishes a new theoretical foundation for pruning algorithm design in LLMs, potentially paving the way for more efficient LLM inference on resource-constrained devices.
@@ 1 Introduction
Large Language Models (LLMs) based on the transformer architecture [136], including GPT-4o [109], Claude [5], and OpenAI’s recent o1 [110], have shown immense potential to enhance our daily lives. They revolutionize fields like conversational AI [78], AI agents [143, 28], search AI [110], and AI assistants [106, 160, 69, 44]. With their growing capabilities, LLMs are powerful tools shaping the future of technology. However, the current state-of-the-art LLM weights number is extremely large. For instance, the smallest version of Llama 3.1 [100] needs 8 billion parameters, which takes more than 16GB GPU memory with half float precision and requires significant inference time. Deploying such models on edge devices such as smartphones becomes challenging due to their large memory and high computational cost. To reduce the LLM model size, many studies work on pruning the LLMs model weights to relax the device memory constraint and minimize response latency. The classical pruning problem in LLMs can be formulated as follows. Given a weight matrix $W\in\mathbb{R}^{d\times d}$ and some calibration data $X\in\mathbb{R}^{n\times d}$, where $n$ is input token length, and $d$ is feature dimension, the goal is to find a matrix $\widetilde{W}$ under some sparse constraint such that $\|XW-X\widetilde{W}\|$ being small under some norm function. The above formulation has been widely used in many state-of-the-art pruning methods, such as SparseGPT [40] and Wanda [120].
However, the current object functions only focus on the approximation of a linear function $XW$. Their optimal solutions do not have a good approximation to the attention matrix (see Figure 2 for details). Note that the attention mechanism is the kernel module of the transformer architecture. The high-level insight of their bad performance is that the $\mathsf{Softmax}$ function is very sensitive to the large positive values of the input due to its $\exp$ scaling effect, while pruning mask based on linear approximation cannot capture this sensitivity. Thus, in this work, we directly compute the pruning mask on weights to approximate the attention matrix, which is a highly non-linear function, $\mathsf{Softmax}(XWX^{\top})\in\mathbb{R}^{n\times n}$. To the best of our knowledge, this paper is the first work studying attention weight pruning to directly approximate the attention matrix. We provide a theoretical guarantee that optimization based on Gradient Descent (GD) on our loss function can converge to a good pruning mask solution (Theorem 1.3). Furthermore, we preliminarily verified the effectiveness of our method with empirical support (Section 6). Our theoretical foundation may pave the way for more efficient LLM inference on resource-constrained devices.
@@
Key background. We introduce some key backgrounds. We define the attention matrix in the self-attention mechanism as below:
@@
Let $M\in[0,1]^{d\times d}$ be the pruning mask. Let $X,W$ be defined in Definition 1.1. Let $A:=\exp(XWX^{\top})$ and $\widetilde{A}:=\exp(X(M\circ W)X^{\top})$, where $\circ$ is the Hadamard product. Let $D:=\operatorname{diag}(A\cdot{\bf 1}_{n})$ and $\widetilde{D}:=\operatorname{diag}(\widetilde{A}\cdot{\bf 1}_{n})$. Let $\lambda\in\mathbb{R}_{+}$ be the regularization parameter. We define the Attention Weights Pruning loss function to be
@@
Our contributions. This is the first work studying the Attention Weights Pruning problem, which is an approximation problem to a non-linear function. We provide an algorithm for obtaining the near-optimal pruning mask based on Gradient Descent (GD) with a convergence guarantee.
@@
In the above theorem, $\xi$ can be arbitrarily small as $\xi\rightarrow 0$ when the regularization coefficient $\lambda\rightarrow 0$. Thus, our analysis shows that although the objective function is highly non-linear, the GD training can converge to a near-optimal pruning mask solution.
Our experiments on synthetic clearly align and support our theoretical analysis (Section 6.1). Furthermore, we evaluate our non-linear pruning method and show it beyond SparseGPT and Wanda by a large margin on real data (C4 Dataset [115]) under real LLM (Llama 3.2-1B [102]) in Section 6.2 and Section 6.3.
@@
  • We conduct experiments to verify the effectiveness of our method (Section 6), showing our non-linear pruning method is beyond SparseGPT and Wanda by a large margin.
@@
In Section 2, we review the related work. Section 3 introduces key concepts and definitions essential for the subsequent sections. In Section 4, we present our main result. Section 5 offers a technical overview of the methods employed. Experimental results are discussed in Section 6. Finally, Section 7 summarizes our findings and offers concluding remarks.
@@ 2.1 Pruning and Compression for LLMs
Model compression plays a critical role in improving the efficiency and deployment of large language models (LLMs) [162] for its effectiveness in reducing computational overhead while preserving performance. Common compression techniques include quantization [112, 146, 55], pruning [19, 51, 52, 66, 39, 40, 120, 88, 156, 168, 153, 2, 25], and knowledge distillation [57, 123, 65, 140]. Specifically, pruning techniques have been developed extensively, such as unstructured pruning, which removes individual weights [88, 120], and structured pruning, which eliminates entire components like neurons or attention heads [104, 2, 144]. [120] proposed Wanda, a novel unstructured pruning technique that uses weight-activation products to induce up to 50% sparsity in LLMs without retraining, achieving competitive results with significantly lower computational cost. SparseGPT [40] introduced a one-shot pruning method that achieves up to 60% sparsity in large GPT-family models with minimal impact on performance. A follow-up work [88] improved the complexity analysis of SparseGPT, reducing the running time from $O(d^{3})$ to $O(d^{2.53})$, enabling faster pruning on LLMs. Together, these techniques contribute to more scalable and resource-efficient LLMs, maintaining competitive performance while substantially reducing computational resources.
@@ 2.2 Attention Acceleration
The attention mechanism has faced criticism due to its quadratic time complexity with respect to context length [136]. Addressing this criticism, a variety of approaches are employed, including sparse attention [52, 70, 40, 88], low-rank approximations [116, 82, 59, 161, 58, 89], and kernel-based methods [20, 101, 34, 158, 94], to reduce computational overhead and improve scalability. [1] enable the derivation of a low-rank representation of the attention matrix, which accelerates both the training and inference processes of single attention layer, tensor attention, and multi-layer transformer, achieving almost linear time complexity [6, 8, 9, 97, 92, 71, 63]. Other approaches like Mamba [45, 32], Linearizing Transformers [155, 105], Hopfield Models [64, 139, 56, 145, 137, 53, 54, 61], and PolySketchFormer [74] focus on architectural modifications and implementation optimizations to enhance performance. System-level optimizations such as FlashAttention [31, 29, 117] and block-wise parallel decoding [124] further improve efficiency. Collectively, these innovations have significantly augmented transformer models’ ability to handle longer input sequences, unlocking broader applications across multiple sectors [27, 114, 38, 107, 150, 4, 67, 84, 95, 121].
@@ 3 Preliminary
Notations. For any positive integer $n$, we use $[n]$ to denote set $\{1,2,\cdots,n\}$. For each $a,b\in\mathbb{R}^{n}$, we use $a\circ b\in\mathbb{R}^{n}$ to denote the Hadamard product, i.e., the $i$-th entry of $(a\circ b)$ is $a_{i}b_{i}$ for all $i\in[n]$. For $A\in\mathbb{R}^{m\times n}$, let $A_{i}\in\mathbb{R}^{n}$ denote the $i$-th row and $A_{*,j}\in\mathbb{R}^{m}$ denote the $j$-th column of $A$, where $i\in[m]$ and $j\in[n]$. We use $\exp(A)$ to denote a matrix where $\exp(A)_{i,j}:=\exp(A_{i,j})$ for a matrix $A\in\mathbb{R}^{n\times d}$. We use $\|A\|_{F}$ to denote the Frobenius norm of a matrix $A\in\mathbb{R}^{n\times d}$, i.e., $\|A\|_{F}:=\sqrt{\sum_{i\in[n]}\sum_{j\in[d]}|A_{i,j}|^{2}}$. For a symmetric matrix $A\in\mathbb{R}^{n\times n}$, $A\succeq 0$ means that $A$ is positive semidefinite (PSD), i.e., for all $x\in\mathbb{R}^{n}$, we have $x^{\top}Ax\geq 0$.
@@
Attention weights pruning. We aim to determine a near-optimal pruning mask $M$ for the attention weights in a self-attention mechanism. Furthermore, we incorporate causal attention masking [^1] In this paper, we always use pruning mask to refer $M\in\mathbb{R}^{d\times d}$, which is our target, and use causal attention mask to refer $M_{c}\in\mathbb{R}^{n\times n}$, which is a fixed mask for standard self-attention. into our method to be more aligned with the current decoder-only LLM architecture, while our analysis can be applied to any general attention mask, e.g., block-wise attention mask. To formalize this, we provide the formal definition of causal attention mask and attention weights pruning in this section.
The causal attention mask ensures that each token in the sequence can attend only to itself and the preceding tokens. Here, we provide the formal definition of the causal attention mask:
Definition 3.1 (Causal attention mask, [92]).
Let $M_{c}\in\{0,1\}^{n\times n}$ be the causal attention mask defined in Definition 3.1. Let $A:=\exp(XWX^{\top})\circ M_{c}$ and $\widetilde{A}:=\exp(X(M\circ W)X^{\top})\circ M_{c}$. Let $D:=\operatorname{diag}(A\cdot{\bf 1}_{n})$ and $\widetilde{D}:=\operatorname{diag}(\widetilde{A}\cdot{\bf 1}_{n})$. We define Attention Weights Pruning with Causal Attention Mask loss function to be $\mathcal{L}(M):=\mathcal{L}_{\mathrm{attn}}(M)+\mathcal{L}_{\mathrm{reg}}(M)$ where $\mathcal{L}_{\mathrm{attn}}(M):=\frac{1}{2}\|D^{-1}A-\widetilde{D}^{-1}\widetilde{A}\|_{F}^{2}~{}~{}\mathrm{and}~{}~{}\mathcal{L}_{\mathrm{reg}}(M):=\frac{1}{2}\lambda\|M\|_{F}^{2}.$
@@ 4 Main Results
We provide an Algorithm 1 for Attention Weights Pruning problem based on Gradient Descent (GD). We also prove the convergence for our GD algorithm in Theorem 4.1.
@@
Let $M,~{}X,~{}W,\widetilde{D},\widetilde{A}~{}\lambda,~{}\mathcal{L},~{}\mathcal{L}_{\mathrm{attn}}$ be defined in Definition 3.2. Assume $XX^{\top}\succeq\beta I$ and $\min_{i,j\in[n]}(\widetilde{D}^{-1}\widetilde{A})_{i,j}\geq\delta>0$. Furthermore, Let $\mu=2\min_{i,j\in[d]}\{|W_{i,j}|\}\cdot\beta\cdot\delta$. Let $\xi=12n^{-1.5}\max_{i,j\in[d]}\{|W_{i,j}|\}\cdot\|X\|_{F}^{2}\cdot\lambda d/\mu$. Then, for any $\epsilon>0$, provided $\eta<1/L$ where $L$ is the Lipschitz constant for $\nabla_{M}\mathcal{L}(M)$ (see Theorem 5.4), GD (Algorithm 1) with fixed step size $\eta$ and run for $T=4\mathcal{L}(M^{(0)})/(\eta\mu\epsilon n^{2})$ iterations results in the following guarantee,
@@
After solving the optimization problem, we obtain a pruning mask with real-valued entries. In practice, however, this pruning mask must be converted into a binary form, specifically $M\in\{0,1\}^{d\times d}$. We define the pruning ratio $\rho$ as the percentage of weights to be pruned. We apply this ratio by setting the pruning mask entries to zero for weights that fall below the $\rho$-th percentile and to one for those above. This ensures that only the specified proportion of weights is pruned.